# Unified People Tracking with Graph Neural Networks

### v1

@@ 4.3 Annotation
We rely on a semi-automated process to annotate our dataset. First, we extract monocular trajectories using a pre-trained version of ByteTrack (Zhang et al., 2022). Next, we convert the detected bounding boxes to 3D bounding cylinders using the calibration and the reconstructed mesh. We then merge 3D trajectories that are close to each other, representing the same object viewed by different cameras. Finally, we developed a tool to visualize and clean up the annotations directly in 3D, displaying all overlapping views at once. Manual cleanup is performed at 1 FPS, and once completed, the annotations are linearly interpolated to 10 FPS.
@@ 5 Experiments
We evaluate our approach on the three datasets of Section 4.
@@ 5.1 Metrics and Implementation Details
We evaluate tracking performance using standard metrics from the literature. For both single-view and multi-view scenarios, we report CLEAR MOT metrics (Kasturi et al., 2009) - MOTA (tracking accuracy) and MOTP (localization precision), along with IDF1 (Ristani et al., 2016) for identity preservation. For single-view tracking, we additionally report HOTA (Luiten et al., 2020). Multi-view metrics are computed in 3D using py-motmetrics, while single-view metrics use trackeval (Luiten & Hoffhues, 2020) in 2D space.
@@
We maintain a rolling window of size $W=10$ frames - once the graph reaches this size, we prune the oldest frame’s nodes and edges when adding new ones, as described in Section 3.1. When pruning, we extract edge and vertex scores which are fed into our trajectory extraction algorithm. We discard vertices and edges with probabilities below thresholds $\tau_{n}=0.5$ and $\tau_{e}=0.5$ respectively.
@@
WILDTRACK KSP-DO [12] 69.6 61.5 73.2
KSP-DO-ptrack [12] 72.2 60.3 78.4
GLMB-YOLOv3 [44] 69.7 73.2 74.3
GLMB-DO [44] 70.1 63.1 72.5
DMCT [65] 72.8 79.1 77.8
DMCT Stack [65] 74.6 78.9 81.9
ReST [13] 84.9 84.1 86.7
EarlyBird [58] 89.5 86.6 92.3
MVFlow [17] 91.3 - 93.5
TrackTacular [59] 91.8 85.4 95.3
@@
Val SORT [8] 75.2 75.2 65.3
ByteTrack [66] 76.8 76.9 66.2
@@
Test CenterTrack [68] 67.8 64.7 52.2
TransTrack [54] 75.2 63.5 54.1
DiffMOT [39] 79.8 79.3 64.5
GeneralTrack [45] 80.6 78.3 64.0
ByteTrack [66] 80.3 77.3 63.1
@@
Test TransTrack [54] 65.0 59.4 48.5
GeneralTrack [45] 77.2 74.0 61.4
ByteTrack [66] 77.8 75.2 61.3
@@
Table 3: Results on MOT17 and MOT20 using the MOTChallenge (Milan et al., 2016) private evaluation protocol. Best results in bold.
@@ 5.2 Benchmarking
We use the WILDTRACK dataset to evaluate our method in a multi-view setting. To represent detection we use their 2D bounding box coordinates, world coordinates, view index and time-step. World coordinates are derived from the 2D bounding box coordinates and the camera calibration using the flat ground plane assumption (Chavdarova et al., 2018). The encoding dimension of those attributes yield vertex feature dimension of $s=704$. We use a pre-trained people detector (Engilberge et al., 2023b) comparable to previous work (Engilberge et al., 2023a; Teepe et al., 2024a).
We report our results in Table 2. Our method significantly outperforms previous state-of-the-art (SOTA) methods boosting MOTA by 2.6% to 93.9. In terms of identity preservation (IDF1), it achieves a nearly 3% improvement over the SOTA reaching 96.3.
@@
For single-view evaluation, we use the MOT17 and MOT20 datasets. Each detection is represented by a feature vector containing its 2D bounding box coordinates, timestamp, and confidence score, yielding a vertex feature dimension of $s=640$ for MOT17 and $s=300$ for MOT20. As in (Zhang et al., 2022; Sun et al., 2020; Qin et al., 2024), we use a pre-trained YOLOX people detector (Ge et al., 2021). Since this paper focuses on linking detections, rather than improving them, this makes for a fair comparison. As can be seen in Tab 3, we outperform ByteTrack on validation on both MOT20 and MOT17. On the private test set, we surpass it in terms of MOTA on MOT17 and match its performance on MOT20, though our IDF1 score lags behind on MOT17. We attribute this performance gap to memory constraints during training, limiting our temporal connections to 6 frames, while ByteTrack can handle gaps of up to 30 frames.
For completeness, we also report the numbers for more recent methods. On MOT17, DiffMOT (Lv et al., 2024) and GeneralTrack (Qin et al., 2024) outperform us on IDF1 and HOTA, but we outperform them on MOTA. This can be explained by the fact that they rely heavily on appearance information to associate detections, which we do not use. On the other hand, appearance seems to be detrimental on the crowded MOT20 where we outperform GeneralTrack on all metrics. In any case, the fact that our numbers are close without using appearance is encouraging and points to an obvious direction for future research: Exploiting appearance features while being able to handle single and multi-view information.
@@
SORT [8] 22.7 - 35.8 33.8
ByteTrack [66] 35.7 - 43.9 49.4
@@
ByteTrack MV 62.1 85.1 62.3
UMPN (Ours) 79.0 86.3 83.2 -
UMPN (Ours) + SP 81.3 87.0 83.3 -
@@
We evaluate on our SCOUT dataset, which provides both single and multi-view tracking scenarios. We use the same detection attributes as in WILDTRACK but we reduce the feature dimensions to $s=224$ in the multi-view case to fit the larger graph in memory. We report our comparative results in Table 4. Overall the metrics are significantly lower than on MOT17 and WILDTRACK because the SCOUT dataset is more demanding than previous ones. Nevertheless, for single-view evaluation, we significantly outperform SORT (Bewley et al., 2016) and ByteTrack (Zhang et al., 2022) and incorporating scene information gives us an additional boost.
In the multi-view setting, existing methods (Engilberge et al., 2023a; Teepe et al., 2024a; b) could not be run because their memory requirements were too high. To nevertheless provide a multi-view baseline, we modified ByteTrack (Zhang et al., 2022) to cluster detections across views and average their world coordinates per cluster before tracking. Instead of 2D IoU, we used 3D IoU between cuboids for similarity computation. Both our models, with and without environmental knowledge, outperform the baseline, with the former achieving slightly better results.
@@ 5.3 Further Analysis
On the WILDTRACK dataset, we test the effect of individual detection attributes used to initialize the vertex features. We test bounding box coordinate, world coordinate, view index, timestep and their combination. You can find results in the top part of Table 6. When using single attributes we see that the bounding box coordinate alone outperforms world coordinate this is consistent with observations from previous work (Bergmann et al., 2019). The combination of the two improves IDF1. Interestingly, combining bbox with view index is beneficial for IDF1. Using the timestep improves performance across all metrics.
@@
Since our model heavily relies on the graph structure, we analyze how different graph construction choices affect tracking performance on MOT17 (Table 5).
@@
Max Temp. dist. $1$ 77.34 59.4 56.5
Max Temp. dist. $4$ 81.4 72.5 64.9
Max Temp. dist. $8$ 81.6 74.9 66.1
@@
We analyze the impact of temporal edge pruning on IDF1 score. Limiting edges to 1 frame reduces performance, as objects cannot be recovered after brief occlusions. Extending to 4 frames (the training value) improves results, with the best performance seen at 6 frames. Increasing to 8 frames, however, proves ineffective, likely due to the training distribution mismatch.
@@
We evaluate key training elements on WILDTRACK, with results in Table 6. Higher embedding dimensionality ($s$) improves performance up to $s=704$, though GPU memory limits further exploration. Larger temporal windows ($W$) boost tracking quality, with $W=10$ improving MOTA by 1.5% compared to $W=5$. Random edge dropout during training enhances robustness, while our delayed backpropagation strategy proves crucial - replacing it with standard backpropagation causes performance to drop by more than 50% on all metrics.
@@
While our method is a learned approach, it maintains competitive computational efficiency compared to handcrafted methods. Timing information is provided in Table 7. The runtime is in the range of other state-of-the-art methods and of the YOLOx detection network. Making it suitable for real-time applications with an average frame rate of 23 FPS.
@@
SORT [8] 17ms - 12ms 29ms 34
ByteTrack [66] 17ms - 13ms 30ms 33
@@ A.1 Graph with Scene Structure
In similar fashion to the detection vertices, the camera vertices are initialized using a two-layer MLP to encode their attributes. The camera vertices encode the camera’s world coordinates $\mathbf{w}_{c}$ along with its intrinsic parameters $\mathbf{K}$ and extrinsic parameters $[\mathbf{R}|\mathbf{t}]$.
@@ A.3 From Probabilistic Graph to Trajectories
The pseudocode of our greedy approach is presented in Algorithm 1. Our algorithm extracts trajectories from the probabilistic graph by merging detections that are connected by high-confidence edges while maintaining spatio-temporal consistency. The key constraint, enforced by the ValidMerge function, ensures that a trajectory cannot contain multiple detections from the same timestamp and camera view, as this would violate physical constraints. The algorithm processes both temporal edges (connecting detections across time) and view edges (connecting simultaneous detections across cameras) to build globally consistent trajectories. By using a disjoint set data structure, we can efficiently merge compatible detections while maintaining the consistency constraints.
@@ A.4 Memory Consumption
We study the GPU memory consumption of our model at training and inference time for different maximum temporal distances between connected vertices. Results can be found in Table 8. The memory consumption at training grows significantly with the increase of the maximum temporal distance. This is expected as the delayed backpropagation requires to store more gradients for any additional eges. At inference time the memory consumption is significantly lower and only grow marginally allowing for edges modeling much longer temporal distances.
@@ A.5 Effect of Temporal Distance
We evaluate the impact of the maximum temporal distance between connected vertices at inference time. As shown in Fig. 5, increasing this distance beyond the training value of 4 initially improves tracking metrics, with peaks at distances 5-6. This suggests that allowing longer-range temporal connections helps recover more challenging trajectories. However, performance eventually degrades at larger distances, likely because the model operates outside its training regime. This highlights a limitation of our current approach - while longer temporal connections are beneficial, memory constraints during training prevent us from directly optimizing for them.
@@ A.6 Optimality of Trajectories Extraction Algorithm
We use a probabilistic objective function to evaluate the optimality of the solutions produced by our algorithm. Following the optimization framework introduced in (6), we compute edge costs as the negative log-likelihood ratios of the edge probabilities. The goal is to minimize the sum of these costs for the edges along the trajectories in a given solution.
@@ A.7 SCOUT Dataset Statistics
SCOUT is also the only dataset that provides a structured 3D representation of the scene as a mesh, offering a richer spatial context for tracking and analysis. Fig. 6 presents example frames from the dataset.
@@
A.8 SCOUT Videos
The supplementary archive includes video visualizations of results on the single-view tracking task. We provide results for the baseline trackers ByteTrack and SORT, alongside the performance of our best model, UMPN+SP.
A.9 Code

### v2

@@ 4.3 Annotation
We rely on a semi-automated process to annotate our dataset. First, we extract monocular trajectories using a pre-trained version of ByteTrack (Zhang22d). Next, we convert the detected bounding boxes to 3D bounding cylinders using the calibration and the reconstructed mesh. We then merge 3D trajectories that are close to each other, representing the same object viewed by different cameras. Finally, we developed a tool to visualize and clean up the annotations directly in 3D, displaying all overlapping views at once. Manual cleanup is performed at 1 FPS, and once completed, the annotations are linearly interpolated to 10 FPS. Due to time limitation, only the first 8 cameras in the first sequence are manually cleaned up. The rest of the dataset is provided with noisy pseudo-labels obtained from the automated part of the annotation pipeline.
@@ 5 Experiments
We evaluate our approach on the three datasets of Section˜4.
@@ 5.1 Metrics and Implementation Details
We evaluate tracking performance using standard metrics from the literature. For both single-view and multi-view scenarios, we report CLEAR MOT metrics (Kasturi09) - MOTA (tracking accuracy) and MOTP (localization precision), along with IDF1 (Ristani16) for identity preservation. For single-view tracking, we additionally report HOTA (Luiten20). Multi-view metrics are computed in 3D using py-motmetrics, while single-view metrics use trackeval (Luiten20b) in 2D space.
@@
We maintain a rolling window of size $W=10$ frames - once the graph reaches this size, we prune the oldest frame’s nodes and edges when adding new ones, as described in Section˜3.1. When pruning, we extract edge and vertex scores which are fed into our trajectory extraction algorithm. We discard vertices and edges with probabilities below thresholds $\tau_{n}=0.5$ and $\tau_{e}=0.5$ respectively.
@@
WILDTRACK KSP-DO [Chavdarova18a] 69.6 61.5 73.2
KSP-DO-ptrack [Chavdarova18a] 72.2 60.3 78.4
GLMB-YOLOv3 [Ong20] 69.7 73.2 74.3
GLMB-DO [Ong20] 70.1 63.1 72.5
DMCT [you20b] 72.8 79.1 77.8
DMCT Stack [you20b] 74.6 78.9 81.9
ReST [Cheng23d] 84.9 84.1 86.7
EarlyBird [Teepe24] 89.5 86.6 92.3
MVFlow [Engilberge23a] 91.3 - 93.5
TrackTacular [Teepe24a] 91.8 85.4 95.3
@@
Val SORT [Bewley16] 75.2 75.2 65.3
ByteTrack [Zhang22d] 76.8 76.9 66.2
@@
Test CenterTrack [Zhou20b] 67.8 64.7 52.2
TransTrack [Sun20b] 75.2 63.5 54.1
DiffMOT [Lv24] 79.8 79.3 64.5
GeneralTrack [Qin24b] 80.6 78.3 64.0
ByteTrack [Zhang22d] 80.3 77.3 63.1
@@
Test TransTrack [Sun20b] 65.0 59.4 48.5
GeneralTrack [Qin24b] 77.2 74.0 61.4
ByteTrack [Zhang22d] 77.8 75.2 61.3
@@
Table 3: Results on MOT17 and MOT20 using the MOTChallenge (Milan16b) private evaluation protocol. Best results in bold.
@@ 5.2 Benchmarking
We use the WILDTRACK dataset to evaluate our method in a multi-view setting. To represent detection we use their 2D bounding box coordinates, world coordinates, view index and time-step. World coordinates are derived from the 2D bounding box coordinates and the camera calibration using the flat ground plane assumption (Chavdarova18a). The encoding dimension of those attributes yield vertex feature dimension of $s=704$. We use a pre-trained people detector (Engilberge23b) comparable to previous work (Engilberge23a; Teepe24).
We report our results in Table˜2. Our method significantly outperforms previous state-of-the-art (SOTA) methods boosting MOTA by 2.6% to 93.9. In terms of identity preservation (IDF1), it achieves a nearly 3% improvement over the SOTA reaching 96.3.
@@
For single-view evaluation, we use the MOT17 and MOT20 datasets. Each detection is represented by a feature vector containing its 2D bounding box coordinates, timestamp, and confidence score, yielding a vertex feature dimension of $s=640$ for MOT17 and $s=300$ for MOT20. As in (Zhang22d; Sun20b; Qin24b), we use a pre-trained YOLOX people detector (Ge21). Since this paper focuses on linking detections, rather than improving them, this makes for a fair comparison. As can be seen in Tab 3, we outperform ByteTrack on validation on both MOT20 and MOT17. On the private test set, we surpass it in terms of MOTA on MOT17 and match its performance on MOT20, though our IDF1 score lags behind on MOT17. We attribute this performance gap to memory constraints during training, limiting our temporal connections to 6 frames, while ByteTrack can handle gaps of up to 30 frames.
For completeness, we also report the numbers for more recent methods. On MOT17, DiffMOT (Lv24) and GeneralTrack (Qin24b) outperform us on IDF1 and HOTA, but we outperform them on MOTA. This can be explained by the fact that they rely heavily on appearance information to associate detections, which we do not use. On the other hand, appearance seems to be detrimental on the crowded MOT20 where we outperform GeneralTrack on all metrics. In any case, the fact that our numbers are close without using appearance is encouraging and points to an obvious direction for future research: Exploiting appearance features while being able to handle single and multi-view information.
@@
SORT [Bewley16] 22.7 - 35.8 33.8
ByteTrack [Zhang22d] 35.7 - 43.9 49.4
@@
ByteTrack MV 19.0 25.7 24.3 -
UMPN (Ours) 24.1 26.1 26.9 -
UMPN (Ours) + SP 25.0 27.8 27.0 -
@@
We evaluate on our SCOUT dataset, which provides both single and multi-view tracking scenarios. We use the same detection attributes as in WILDTRACK but we reduce the feature dimensions to $s=224$ in the multi-view case to fit the larger graph in memory. We split the manually cleaned dataset (first 8 cameras in the first sequence) into two equal parts: the first is used for training, and the second for evaluation. We report our comparative results in Table˜4. Overall the metrics are significantly lower than on MOT17 and WILDTRACK because the SCOUT dataset is more demanding than previous ones. Nevertheless, for single-view evaluation, we significantly outperform SORT (Bewley16) and ByteTrack (Zhang22d) and incorporating scene information gives us an additional boost.
In the multi-view setting, existing methods (Engilberge23a; Teepe24; Teepe24a) could not be run because their memory requirements were too high. To nevertheless provide a multi-view baseline, we modified ByteTrack (Zhang22d) to cluster detections across views and average their world coordinates per cluster before tracking. Instead of 2D IoU, we used 3D IoU between cuboids for similarity computation. Both our models, with and without environmental knowledge, outperform the baseline, with the former achieving slightly better results.
@@ 5.3 Further Analysis
On the WILDTRACK dataset, we test the effect of individual detection attributes used to initialize the vertex features. We test bounding box coordinate, world coordinate, view index, timestep and their combination. You can find results in the top part of Table˜6. When using single attributes we see that the bounding box coordinate alone outperforms world coordinate this is consistent with observations from previous work (Bergmann19a). The combination of the two improves IDF1. Interestingly, combining bbox with view index is beneficial for IDF1. Using the timestep improves performance across all metrics.
@@
Since our model heavily relies on the graph structure, we analyze how different graph construction choices affect tracking performance on MOT17 (Table˜5).
@@
Max Temp. dist. 11 77.34 59.4 56.5
Max Temp. dist. 44 81.4 72.5 64.9
Max Temp. dist. 88 81.6 74.9 66.1
@@
We analyze the impact of temporal edge pruning on IDF1 score. At inference time, limiting edges to 1 frame reduces performance, as objects cannot be recovered after brief occlusions. Extending to 4 frames (the training value) improves results, with the best performance seen at 6 frames. Increasing to 8 frames, however, proves ineffective, likely due to the training distribution mismatch.
@@
We evaluate key training elements on WILDTRACK, with results in Table˜6. Higher embedding dimensionality ($s$) improves performance up to $s=704$, though GPU memory limits further exploration. Larger temporal windows ($W$) boost tracking quality, with $W=10$ improving MOTA by 1.5% compared to $W=5$. Random edge dropout during training enhances robustness, while our delayed backpropagation strategy proves crucial - replacing it with standard backpropagation causes performance to drop by more than 50% on all metrics.
@@
While our method is a learned approach, it maintains competitive computational efficiency compared to handcrafted methods. Timing information is provided in Table˜7. The runtime is in the range of other state-of-the-art methods and of the YOLOx detection network. Making it suitable for real-time applications with an average frame rate of 23 FPS.
@@
SORT [Bewley16] 17ms - 12ms 29ms 34
ByteTrack [Zhang22d] 17ms - 13ms 30ms 33
@@ A.1 Graph with Scene Structure
In similar fashion to the detection vertices, the camera vertices are initialized using a two-layer MLP to encode their attributes. The camera vertices encode the camera’s world coordinates $\mathbf{w}_{c}$ along with its intrinsic parameters $\mathbf{K}$ and extrinsic parameters $\mathbf{R}|\mathbf{t}]$.
@@ A.3 From Probabilistic Graph to Trajectories
The pseudocode of our greedy approach is presented in Algorithm˜1. Our algorithm extracts trajectories from the probabilistic graph by merging detections that are connected by high-confidence edges while maintaining spatio-temporal consistency. The key constraint, enforced by the ValidMerge function, ensures that a trajectory cannot contain multiple detections from the same timestamp and camera view, as this would violate physical constraints. The algorithm processes both temporal edges (connecting detections across time) and view edges (connecting simultaneous detections across cameras) to build globally consistent trajectories. By using a disjoint set data structure, we can efficiently merge compatible detections while maintaining the consistency constraints.
@@ A.4 Memory Consumption
We study the GPU memory consumption of our model at training and inference time for different maximum temporal distances between connected vertices. Results can be found in Table˜8. The memory consumption at training grows significantly with the increase of the maximum temporal distance. This is expected as the delayed backpropagation requires to store more gradients for any additional eges. At inference time the memory consumption is significantly lower and only grow marginally allowing for edges modeling much longer temporal distances.
@@ A.5 Effect of Temporal Distance
We evaluate the impact of the maximum temporal distance between connected vertices at inference time. As shown in Fig.˜5, increasing this distance beyond the training value of 4 initially improves tracking metrics, with peaks at distances 5-6. This suggests that allowing longer-range temporal connections helps recover more challenging trajectories. However, performance eventually degrades at larger distances, likely because the model operates outside its training regime. This highlights a limitation of our current approach - while longer temporal connections are beneficial, memory constraints during training prevent us from directly optimizing for them.
@@ A.6 Optimality of Trajectories Extraction Algorithm
We use a probabilistic objective function to evaluate the optimality of the solutions produced by our algorithm. Following the optimization framework introduced in (Berclaz11), we compute edge costs as the negative log-likelihood ratios of the edge probabilities. The goal is to minimize the sum of these costs for the edges along the trajectories in a given solution.
@@ A.7 SCOUT Dataset Statistics
SCOUT is also the only dataset that provides a structured 3D representation of the scene as a mesh, offering a richer spatial context for tracking and analysis. Fig.˜6 presents example frames from the dataset.
@@
Figure 7: Scene occlusion challenges in SCOUT dataset. Two sequences of 4 frames each illustrate the challenging occlusion scenarios present in our dataset. People are marked with colored dots; when occluded, the dots display a checkerboard pattern. Top sequence: Basic occlusion scenarios where individuals become temporarily occluded by scene structures. Bottom sequence: More challenging repetitive scene occlusion as a group of people walks along columns, creating systematic and recurring occlusion patterns that are particularly difficult for tracking algorithms to handle.
A.8 Code